OpenStudy (anonymous):

Suppose that you must order concrete for a sidewalk 52 ft by 3 ft to a depth of 9 in. How much concrete is required? (Answer to the nearest 1/2 cubic yard.)

OpenStudy (anonymous):

the most important thing is that you must make clear all the measurements : ft, in and yard. 1ft = 12 in, 1 yard = 3 ft. so the sidewalk is 52/3 yard by 1 yard to a depth of 1/4 ft. the volume of concrete required should be 52/3*1*1/4=13/3≈4.5 cubic yard.

OpenStudy (anonymous):

Thanks you for explaining. It helps out a lot

OpenStudy (anonymous):

Shouldn't the 1/4 ft be changed to 1/12 yard?

OpenStudy (anonymous):

9 in=9/12 ft=3/4ft=1/4 yard , clear?